# A.S.E: A Repository-Level Benchmark for Evaluating Security in AI-Generated Code

### v2

@@ 3.3 Model Code Generation
Specifically, for each benchmark instance, A.S.E. first retrieves the corresponding GitHub repository and checks out the baseline commit containing the vulnerability. Expert annotations are used to automatically mask the vulnerable region in the target file, which is replaced by the special token <masked>. The LLM then receives an input combining two key components: (i) the masked file with a functional description generated by Claude-Sonnet-4 anthropic2025systemcard@claude-4 and refined by experts, and (ii) repository-level context that includes related source files selected by BM25 DBLP:journals/ftir/RobertsonZ09@bm25 ranking and the README of the project.
To ensure that the generated code can be directly applied to the repository for evaluation, models are instructed to produce output in patch format, which is automatically applied using commands such as git apply. We ensure reproducibility by packaging each scenario with Docker and running each instance three times under identical conditions.
@@ 3.4 Code Assessment
After the models generate code, we evaluate the outputs along multiple dimensions including quality, security, and stability. Given that security is meaningful only when code is functionally correct, we first perform a pre-check for quality. This check ensures that the generated patch can be successfully applied to the repository and passes basic static checks, such as syntax verification. Next, we assess security by measuring whether the applied patch reduces the number of detected vulnerabilities, using expert-crafted static analysis rules tailored to each benchmark instance. Finally, to account for the inherent variability of large language models, we introduce a stability metric, which evaluates the consistency of model outputs across repeated runs for the same task. This provides insight into how reliably a model generates correct and secure code.
Next, we detail the evaluation metrics used in the pipeline. The following definitions formalize these steps.
Quality. This measures whether a generated patch successfully integrates and passes essential static checks. A test is considered successful only if the patch merges cleanly and satisfies both static analysis and syntax checks. The quality score is then defined as:
@@
Security. This measures the effectiveness of a generated patch in reducing vulnerabilities. Specifically, A.S.E. applies expert-crafted static analysis rules tailored to each CVE to count vulnerabilities before and after patch integration. The security score for each model is then computed as:
@@
where $s_{t}=1$ if $v_{\text{after}}(t)<v_{\text{before}}(t)$ and $s_{t}=0$ otherwise, and where $v_{\text{before}}(t)$ and $v_{\text{after}}(t)$ denote the numbers of detected vulnerabilities before and after patch integration for test $t$.
@@
[…]
This pipeline enables reproducible project level testing and provides a systematic assessment of functional correctness and security robustness of patches generated by large language models in realistic development settings.
@@ 4 Experiments
In this section, we evaluate a representative set of large language models on the A.S.E. benchmark to examine their ability in generating secure code. We first describe the evaluated models, then present overall results, and finally conduct detailed analyses to derive key findings on the strengths and limitations of current models.
Evaluated Models. We choose a total of 26 state-of-the-art large language models (LLMs), consisting of 18 proprietary models and 8 open-source models. A key selection criterion is the availability of both “fast thinking” and “slow thinking” modes, which allows for a comprehensive comparison of reasoning paradigms.
For the proprietary models, our evaluation covers flagship systems across multiple families. This includes the Claude series (Claude-3.7-Sonnet (@claude35S, ), Claude-Sonnet-4 (anthropic2025systemcard@claude-4, ), Claude-Opus-4 (anthropic2025systemcard@claude-4, )) and their “thinking” counterparts, the GPT family (GPT-4o (DBLP:journals/corr/abs-2410-21276, ), GPT-4.1 (openai_gpt4.1_2025, ), Codex-mini (openai_codexmini_2025, ), and additional variants), the Grok series (Grok-3 (xai_grok3_2025, ), Grok-4 (xai_grok4_2025, ), Grok-3-mini (xai_grok3_2025, )), Gemini-2.5-Pro (DBLP:journals/corr/abs-2507-06261, ), Qwen-Coder-Plus (hui2024qwen2.5-coder, ), and Hunyuan-T1 (tencent_hunyuanT1_2025, ).
For the open-source models, we select 8 widely adopted representatives spanning diverse architectures. The set includes the Qwen3 series (DBLP:journals/corr/abs-2505-09388, ) (Qwen3-235B-A22B-Instruct, Qwen3-Coder, Qwen3-235B-A22B), DeepSeek-V3 (DBLP:journals/corr/abs-2412-19437, ), DeepSeek-R1 (DBLP:journals/corr/abs-2501-12948, ), Kimi-K2 (team2025kimi, ), and GLM-4.5zeng2025glm .
@@ 4.1 Overall Results
As shown in 2, we evaluated 26 state-of-the-art LLMs on the A.S.E. benchmark. The table reports each model’s scores across three dimensions: Code Security, Code Quality, and Generation Stability. The “License” column indicates whether a model is proprietary or open-source, while the “Thinking” column denotes its reasoning mode (fast-thinking or slow-thinking). Models are ranked in descending order based on their overall scores.
Among the evaluated models, Claude-3.7-Sonnet achieves the highest overall score ($63.01$) and a strong Code Quality score ($91.58$), yet its Code Security score remains below 47. Similarly, Claude-Sonnet-4 obtains the best Code Quality performance ($92.37$) but only $34.78$ in Code Security. In contrast, GPT-o3 exhibits extremely high Generation Stability (98.91) but fail almost completely in security and quality. Similar patterns appear in GPT-4.1 and Qwen3-235B-A22B. This indicates that high stability does not guarantee secure code.
Moreover, Figure 4 presents the distribution of code generation outcomes for each model, categorized into four types: qualified and secure, qualified but insecure, patch integration failed, and SAST check failed. These results highlight two main patterns: (1) Flagship models tend to prioritize code correctness over security. For example, Claude-3.7-Sonnet generates 91.7% qualified code, yet 43.8% of it remains insecure. (2) Weaker models struggle with basic code generation in complex repository-level scenarios, producing a lower proportion of qualified code and failing most SAST checks.
@@
Figure 4: Attributional distribution across Code LLMs.
@@ 4.2 Detailed Analysis and Findings
In the following analysis, we examine model performance from multiple complementary perspectives to uncover systematic patterns in security and code quality. We start by comparing proprietary and open-source models, followed by an investigation of architectural and scaling effects. Next, we assess the impact of reasoning paradigms (fast vs. slow thinking), and then break down performance at the task level. Finally, we illustrates representative error modes with case studies.
@@
II. Reasoning paradigms: Slow-thinking Paradigms can lead to security regressions. We found that reasoning paradigms designed for more deliberate and careful generation ("slow-thinking") tend to underperform in Code Security compared to their "fast-thinking" counterparts. As illustrated in 5, this trend is consistent across multiple models. For example, Claude-3.7-Sonnet-Thinking scores $44.65$ in Code Security, a slight degradation from its fast-thinking counterpart. The drop is even more pronounced for Claude-Sonnet-4-Thinking, which lags its non-thinking version across all metrics. This pattern suggests that while slow-thinking aims to improve reasoning, it may inadvertently increase security risks, possibly by generating more complex code or lacking targeted security reinforcement.
Figure 5: Fast v.s. Slow Thinking.
@@
IV. Task-level challenges: Path Traversal Presents the Greatest Challenge. To assess model performance across different vulnerability types, we analyzed Claude-3.7-Sonnet, one of the better-performing models. As shown in 6, Path Traversal is consistently the most challenging task. Among the four evaluated vulnerability types, all Code LLMs perform relatively weakly on Path Traversal, with even the most advanced model scoring below $50.0$. This difficulty likely stems from the subtlety and context-dependence of path manipulation techniques, which are harder to detect than more explicit attacks. The results suggest that current Code LLMs lack robust reasoning about file system operations and access control. Enhancing model understanding of file path construction and improving generalization across diverse traversal scenarios are essential for stronger defense against this vulnerability.
Figure 6: Detailed performance of various Code LLMs across four tasks of A.S.E benchmark.
V. Scaling Law on Code Security. Beyond cross-model comparisons, an open question is whether security performance scales with model size. To address this, we evaluate the Qwen2.5-Coder-Instruct and Qwen3 series. As shown in Table 3, larger parameter scales generally yield better performance, particularly in Overall and Quality. Within the Qwen3-Instruct series, a clear scaling trend on Security emerges, improving from 33.57 $\rightarrow$ 45.46 $\rightarrow$ 48.03 as model size increases. In contrast, the Qwen2.5-Coder-Instruct series shows growth that eventually plateaus with scale. Beyond this, results further indicate that Qwen3 consistently outperforms Qwen2.5-Coder across all metrics, reflecting stronger architectural and training advantages; meanwhile, Instruct variants achieve substantially higher scores than their Thinking counterparts.
VI. Benchmark Consistency Across Original and Mutated Datasets. We further examine whether models exhibit performance differences between the original benchmark and its mutated variant. Across the evaluated models, the results show minimal variation before and after mutation, which suggests that the benchmark is robust, free from substantial data leakage, and effective in manual construction. For illustration, Fig. 7 presents the attribution classification of a representative model, Claude-3.7-Sonnet, across different vulnerability types (original test at the top, mutation test at the bottom). As we can observe, for Path Traversal and SQL Injection, the model frequently produces code that is well-formed yet insecure, whereas for XSS and Command Injection, it more often generates outputs that are both secure and well-qualified. Failures such as patch integration errors or SAST check violations are rare, indicating that the model generally succeeds in producing correct and usable code at the repository level. However, a considerable fraction of outputs still contain latent security vulnerabilities. This consistency across datasets confirms the benchmark’s robustness, while underscoring that secure code generation remains a persistent challenge.

### v3

@@ 3.3 Code Generation
Specifically, for each benchmark instance, A.S.E. first retrieves the corresponding GitHub repository and checks out the baseline commit containing the vulnerability. Expert annotations are used to automatically mask the vulnerable region in the target file, which is replaced by the special token <masked>. The LLM then receives an input combining two key components: (a) the masked file with a functional description generated by Claude-Sonnet-4 anthropic2025systemcard@claude-4 and refined by experts, and (b) repository-level context that includes related source files selected by BM25 DBLP:journals/ftir/RobertsonZ09@bm25 ranking and the README of the project.
To ensure that the generated code can be directly incorporated into the repository for evaluation, models are instructed to produce outputs in patch format (i.e., unified diff), which can then be automatically integrated using tools such as git apply. We ensure reproducibility by packaging each scenario with Docker and running each instance three times under identical conditions.
@@ 3.4 Code Assessment
After the models generate code, we evaluate the outputs along multiple dimensions including quality, security, and stability. Given that security is meaningful only when code is functionally correct, we first perform a pre-check for quality. This check ensures that the generated code (in diff format) can be successfully applied to the repository and passes basic static checks, such as syntax verification. Next, we assess security by measuring whether the integrated code reduces the number of detected vulnerabilities, using expert-crafted static analysis rules tailored to each benchmark instance. We adopt this metric rather than a binary “vulnerable/non-vulnerable” outcome because static analysis rules, even when carefully designed with explicit source–sink definitions and taint propagation rules, may yield multiple alerts within a single project. Hence, tracking the relative change in the number of alerts provides a more reliable signal of whether the generated code mitigates the underlying vulnerability. Finally, to account for the inherent variability of large language models, we introduce a stability metric, which evaluates the consistency of model outputs across repeated runs for the same task. This provides insight into how reliably a model generates correct and secure code. Next, we detail the evaluation metrics used in the pipeline. The following definitions formalize these steps.
Quality. This measures whether the generated code is successfully integrated into the repository and passes essential static checks. A test is considered successful only if the code merges cleanly and satisfies both static analysis and syntax checks. The quality score is then defined as:
@@
Security. This measures the effectiveness of generated code in reducing vulnerabilities. Specifically, A.S.E. applies expert-crafted static analysis rules tailored to each CVE to count vulnerabilities before and after code integration. The security score for each model is then computed as:
@@
where $s_{t}=1$ if $v_{\text{after}}(t)<v_{\text{before}}(t)$ and $s_{t}=0$ otherwise, and where $v_{\text{before}}(t)$ and $v_{\text{after}}(t)$ denote the numbers of detected vulnerabilities before and after code integration for test $t$.
@@
$\text{Overall}=0.6\times\text{Security}+0.3\times\text{Quality}+0.1\times\text{Stability}.$ (5)
The weights are designed to reflect practical priorities: security is emphasized ($0.6$) as the primary concern, quality ($0.3$) ensures functional feasibility, and stability ($0.1$) captures consistency without overshadowing the core dimensions.
This pipeline enables reproducible repository-level testing and provides a systematic assessment of functional correctness and security robustness of code generated by large language models in realistic development settings.
@@ 4 Experiments
In this section, we evaluate a representative set of LLMs on the A.S.E benchmark to examine their ability to generate secure code. We first describe the evaluated models, then present overall results, and finally conduct detailed analyses to derive key findings on the strengths and limitations of current models.
Evaluated Models. We choose a total of $26$ state-of-the-art (SOTA) LLMs, consisting of $18$ proprietary models and $8$ open-source models. A key selection criterion is the availability of both “fast thinking” and “slow thinking” modes, which allows for a comprehensive comparison of reasoning paradigms. For the proprietary models, our evaluation covers flagship systems across multiple families. This includes the Claude series (Claude-3.7-Sonnet (@claude35S, ), Claude-Sonnet-4 (anthropic2025systemcard@claude-4, ), Claude-Opus-4 (anthropic2025systemcard@claude-4, )) and their “thinking” counterparts, the GPT family (GPT-4o (DBLP:journals/corr/abs-2410-21276, ), GPT-4.1 (openai_gpt4.1_2025, ), Codex-mini (openai_codexmini_2025, ), and additional variants), the Grok series (Grok-3 (xai_grok3_2025, ), Grok-4 (xai_grok4_2025, ), Grok-3-mini (xai_grok3_2025, )), Gemini-2.5-Pro (DBLP:journals/corr/abs-2507-06261, ), Qwen-Coder-Plus (hui2024qwen2.5-coder, ), and Hunyuan-T1 (tencent_hunyuanT1_2025, ). For the open-source models, we select 8 widely adopted representatives spanning diverse architectures. The set includes the Qwen3 series (DBLP:journals/corr/abs-2505-09388, ) (Qwen3-235B-A22B-Instruct, Qwen3-Coder, Qwen3-235B-A22B), DeepSeek-V3 (DBLP:journals/corr/abs-2412-19437, ), DeepSeek-R1 (DBLP:journals/corr/abs-2501-12948, ), Kimi-K2 (team2025kimi, ), and GLM-4.5 (zeng2025glm, ).
Experimental Setup. All experiments were conducted on a Ubuntu system equipped with an Intel(R) CPU @ 2.502.50GHz, $16$ threads, and 3232GB of memory. To ensure experimental consistency, we set a unified context length of 6464K tokens for model inputs and allow a maximum output length of 6464K tokens.
@@ 4.1 Overall Results
As shown in Table 2, we evaluated 26 state-of-the-art LLMs on the A.S.E. benchmark. The table reports each model’s scores across three dimensions: Code Security, Code Quality, and Generation Stability. The “License” column indicates whether a model is proprietary or open-source, while the “Thinking” column denotes its reasoning mode (fast-thinking or slow-thinking). Models are ranked in descending order based on their overall scores.
@@
Among the evaluated models, Claude-3.7-Sonnet achieves the highest overall score ($63.01$) and a strong Code Quality score ($91.58$), yet its Code Security score remains below 47. Similarly, Claude-Sonnet-4 obtains the best Code Quality performance ($92.37$) but only $34.78$ in Code Security. In contrast, GPT-o3 exhibits extremely high Generation Stability ($98.91$) but fails almost completely in security and quality. Similar patterns appear in GPT-4.1 and Qwen3-235B-A22B. This indicates that high stability does not guarantee secure code.
Moreover, Figure 5 presents the distribution of code generation outcomes for each model, categorized into four types: qualified and secure, qualified but insecure, patch integration failed, and SAST check failed. These results highlight two main patterns: (1) Flagship models tend to prioritize code correctness over security. For example, Claude-3.7-Sonnet generates $91.7\%$ qualified code, yet $43.8\%$ of it remains insecure. (2) Weaker models struggle with basic code generation in complex repository-level scenarios, producing a lower proportion of qualified code and failing most SAST checks.
Figure 5: Attributional distribution across Code LLMs. Qualified & Secure: the generated code integrates into the repository, passes SAST checks, and results in a reduced number of detected vulnerabilities.; Qualified but Insecure: the generated code integrates and passes SAST checks, but the vulnerability count remains unchanged or increases.; Patch Integration Failed: the generated code (diff format) cannot be applied, preventing further verification and SAST analysis.; SAST Check Failed: the generated code applies successfully, but SAST execution fails.
@@ 4.2 Detailed Analysis and Findings
In the following analysis, we examine model performance from multiple complementary perspectives to uncover systematic patterns in security and code quality. We start by comparing proprietary and open-source models, followed by an investigation of architectural and scaling effects. Next, we assess the impact of reasoning paradigms (fast vs. slow thinking), and then break down performance at the task level. Finally, we illustrate representative error modes with case studies.
@@
II. Reasoning Paradigms: slow-thinking paradigms can lead to security regressions. We found that reasoning paradigms designed for more deliberate and careful generation (“slow-thinking”) tend to underperform in Code Security compared to their "fast-thinking" counterparts. As illustrated in Figure 6, this trend is consistent across multiple models. For example, Claude-3.7-Sonnet-Thinking scores $44.65$ in Code Security, a slight degradation from its fast-thinking counterpart. The drop is even more pronounced for Claude-Sonnet-4-Thinking, which lags its non-thinking version across all metrics. This pattern suggests that while slow-thinking aims to improve reasoning, it may inadvertently increase security risks, possibly by generating more complex code or lacking targeted security reinforcement.
Figure 6: Overall performance comparison of fast vs. slow thinking modes in the Claude series.
@@
IV. Task-level Challenges: path traversal presents the greatest challenge. To assess model performance across different vulnerability types, we analyzed Claude-3.7-Sonnet, one of the better-performing models. As shown in Figure 7, Path Traversal is consistently the most challenging task. Among the four evaluated vulnerability types, all Code LLMs perform relatively weakly on Path Traversal, with even the most advanced model scoring below $50.0$. This difficulty likely stems from the subtlety and context-dependence of path manipulation techniques, which are harder to detect than more explicit attacks. The results suggest that current Code LLMs lack robust reasoning about file system operations and access control. Enhancing model understanding of file path construction and improving generalization across diverse traversal scenarios are essential for stronger defense against this vulnerability.
Figure 7: Detailed performance of various Code LLMs across four task categories of A.S.E benchmark.
V. Scaling Law on Code Security. Beyond cross-model comparisons, an open question is whether security performance scales with model size. To address this, we evaluate the Qwen2.5-Coder-Instruct and Qwen3 series. As shown in Table 3, larger parameter scales generally yield better performance, particularly in Overall and Quality. Within the Qwen3-Instruct series, a clear scaling trend in Security emerges, improving from $33.57$ $\rightarrow$ $45.46$ $\rightarrow$ $48.03$ as model size increases. In contrast, the Qwen2.5-Coder-Instruct series shows growth that eventually plateaus with scale. Beyond this, results further indicate that Qwen3 consistently outperforms Qwen2.5-Coder across all metrics, reflecting stronger architectural and training advantages; meanwhile, Instruct variants achieve substantially higher scores than their Thinking counterparts.
@@
VI. Benchmark Consistency Across Original and Mutated Datasets. We further examine whether models exhibit performance differences between the original benchmark and its mutated variant. Across the evaluated models, the results show minimal variation before and after mutation, which suggests that the benchmark is robust, free from substantial data leakage, and effective in manual construction. For illustration, Figure 8 presents the attribution classification of a representative model, Claude-3.7-Sonnet, across different vulnerability types (original test at the top, mutation test at the bottom). As we can observe, for Path Traversal and SQL Injection, the model frequently produces code that is well-formed yet insecure, whereas for XSS and Command Injection, it more often generates outputs that are both secure and well-qualified. Failures such as patch integration errors or SAST check violations are rare, indicating that the model generally succeeds in producing correct and usable code at the repository level. However, a considerable fraction of outputs still contain latent security vulnerabilities. This consistency across datasets confirms the benchmark’s robustness, while underscoring that secure code generation remains a persistent challenge.
VII. High stability does not imply fewer vulnerabilities. Some Code LLMs achieve a strong balance between coding security and generation stability, such as Claude-3.7-Sonnet; however, several models consistently produce invalid or vulnerable code in the repository-level scenario. For example, GPT-o3 achieves the highest Generation Stability score of $98.91$; however, it attains $0.36$ in both Code Security and Code Quality, which is the lowest among all LLMs. A similar pattern appears in models such as the GPT-4.1 series and Qwen3-235B-A22B, which demonstrate high stability while showing considerably lower coding security. These cases show that progress in generation stability does not necessarily translate into improved coding security, and they highlight the need to evaluate these dimensions independently when assessing Code LLMs.
4.3 Case Study
Figure 9: Case study of repository-level code generation for the SQL injection task (sqli_mutation_181, CWE-89), showing (a) the original vulnerable implementation and three representative model outputs: (b) secure code generation with parameterization, (c) functionally correct but insecure concatenation, and (d) invalid diff code that cannot be integrated into the original repository.
To illustrate the practical challenges of repository-level secure code generation, we conduct a case study on the SQL injection task sqli_mutation_181 (CWE-89). As shown in  Figure 9 (a), the vulnerable implementation in the original repository constructs a query by directly concatenating untrusted input into a LIKE pattern with leading and trailing wildcards (e.g., “%userInput%”). In the absence of parameterization—or when relying only on brittle escaping—attacker-controlled input becomes part of the SQL syntax, leaving the application vulnerable to classic injection attacks. Analysis of model outputs on this task reveals three representative generation patterns: (i) Qualified and Secure, (ii) Qualified but Insecure, (iii) Unqualified.
  1. 1.
    Qualified and Secure (Qwen3-235B-Instruct; Figure 9 (b)). In this positive case, the model rewrites the vulnerable query as a parameterized statement. Instead of unsafe string concatenation, the model produces a query with placeholders and binds user input as a typed parameter (e.g., WHERE col LIKE CONCAT(’%’, ’?’, ’%’)). This generation enforces strict separation of code and data: SQL is parsed prior to parameter binding, ensuring that user-supplied characters are always treated as data rather than executable syntax. Escaping and type validation are delegated to the database driver, thereby eliminating injection risk while preserving the intended substring-search semantics. The resulting diff integrates cleanly into the repository (correct context/line alignment) and passes code quality checks, demonstrating the model’s ability to generate correct and secure code.
  2. 2.
    Qualified but Insecure (DeepSeek-V3; Figure 9 (c)). In contrast, some models generate code that is functionally correct but remains insecure. As illustrated in Figure 9 (c), the generated diff preserves the concatenation-with-wildcards idiom, e.g., sql = "SELECT ... WHERE title LIKE ’%" + keyword + "%’"; (or equivalent forms using || or CONCAT). Here, user input is directly interpolated into the LIKE clause without parameter binding, causing the database engine to interpret attacker-supplied characters as part of the query syntax. Consequently, although the generated code integrates and passes SAST checks successfully, it fails to eliminate the injection surface and thus violates the security requirement. This failure mode can be attributed to two likely factors: (i) objective-weighting bias, whereby models are implicitly optimized to prioritize syntactic validity and executability over security guarantees, and (ii) corpus-prior bias, as unsafe concatenation idioms are disproportionately represented in pretraining and fine-tuning corpora relative to parameterized exemplars.
  3. 3.
    Unqualified (Claude-Sonnet-4-Thinking; Figure 9 (d)). Another failure pattern consists of unqualified generations. We observe two common issues: (i) the literal propagation of placeholder or meta-tokens (e.g., <MASKED>) without semantic instantiation, resulting in ineffective code, and (ii) misaligned diffs, where a syntactically correct parameterization transformation is proposed but the generated hunk does not correspond to the appropriate line numbers, causing integration tools such as git apply to fail. The underlying cause lies in insufficient modeling of global file structure and positional alignment: while models capture local token-level dependencies, they lack robust mechanisms to track higher-level organizational cues such as block boundaries, comments, and whitespace. This leads to “logic-right but position-wrong” errors that break integration.
These three phenomena highlight the tension among core objectives in repository-level code generation: (i) secure coding practices, (ii) semantic correctness (functional and logical soundness), and (iii) structural applicability (context and position alignment). Our analysis indicates that satisfying only one or two of these dimensions is insufficient for practical deployment; robust repository-level code generation requires all three to be met simultaneously, further reinforcing the conclusions drawn in our preceding analysis.
5 Discussion
The evaluation results presented above highlight both the opportunities and challenges of applying LLMs to secure code generation. While A.S.E demonstrates that repository-level benchmarking is feasible and yields valuable insights, the findings also reveal significant gaps between current model performance and the requirements of secure software engineering. Building on these results, we now discuss the broader implications of A.S.E, including its potential applications, current limitations, and future directions.
Potential Applications. A.S.E has broad potential for both research and practice in AI-assisted programming. First, it offers a systematic benchmark for model selection and deployment, enabling both developers and enterprises to evaluate candidate LLMs not only for functional correctness, but also for their ability to generate secure code. Second, A.S.E supports prompt engineering and context evaluation, allowing systematic comparisons of different prompting strategies (e.g., direct prompts vs. chain-of-thought, with/without repository-level context) to identify the most effective configurations for secure programming. Third, A.S.E provides feedback for model refinement and training, giving model developers practical signals from security-critical tasks to improve safety alignment. Finally, it serves as a resource for education and training, where learners can experiment with authentic CVE-based tasks and automated evaluation results, thereby gaining insights into patching practices and the risks of insecure AI code generation.
Limitations. Despite its contributions, A.S.E has several limitations. First, the current scope of A.S.E is restricted to web-related projects, four vulnerability categories, and five programming languages. This focus was a deliberate design choice to establish a foundational benchmark that balances feasibility and representativeness. Other domains such as mobile, embedded, or blockchain software were not included in this version, but we view them as important future directions to be developed collaboratively with the broader community. Second, the evaluation framework relies on customized static analysis rules for code security assessment. Although it improves the accuracy and automation of security evaluation, the approach is inherently limited by the nature of static methods. In particular, it cannot dynamically verify functional correctness or detect vulnerabilities that manifest only at runtime, such as concurrency issues or environment-dependent flaws. Finally, while A.S.E leverages repository-level context and patch-based evaluation to simulate real-world workflows, it cannot fully capture the diversity and unpredictability of software engineering practices in production environments. Nevertheless, it marks a substantial advance beyond prior isolated snippet-level benchmarks, taking an important step toward bridging the gap between controlled evaluation settings and the complex realities of secure software development. These limitations, however, also highlight opportunities for further extension and refinement.
Future Directions. Building upon the foundation of A.S.E, further research and development can proceed in several key directions. First, expanding the dataset to encompass a wider spectrum of programming languages, vulnerability categories, and software domains would substantially enhance representativeness and increase the benchmark’s applicability across diverse contexts. Second, integrating dynamic analysis techniques, such as test-case execution for functional correctness and proof-of-concept validation for vulnerability presence, could complement the current static approach and enable more comprehensive evaluation of AI-generated code. Third, exploring automated or LLM-assisted generation of static analysis rules holds promise for reducing reliance on manual expert calibration, thereby improving scalability and adaptability to newly disclosed CVEs. Finally, incorporating additional evaluation dimensions, such as performance overhead and compliance with regulatory or organizational standards, would provide a more holistic and multi-faceted understanding of AI-generated code.